### v1

@@ Abstract
Keywords:
Quasi-incompressible Navier–Stokes equations Maxwell–Stefan equations cross-diffusion systems finite element method structure-preserving numerical scheme discrete mass conservation energy dissipation
MSC:
65M60 65N30 76T30 80M10.
@@ 1 Introduction
Many applications for multicomponent mixtures, like the ion transport in biological channels, the dynamics of electrolytes in lithium-ion batteries, or helium–oxygen–carbon dioxide mixtures in the lung, can be described by Maxwell–Stefan equations. In the Fick–Onsager form, they consist of nonlinear parabolic cross-diffusion equations whose diffusion matrix (called mobility matrix) is positive semidefinite only BoDr23 . In the hydrostatic and isothermal case, the variables are the partial mass densities that satisfy a volume-filling constraint implying a constant total mass density. In the general case, the equations are coupled to the compressible or incompressible Navier–Stokes equations. In this paper, we consider the quasi-incompressible model of Dru21 and suggest, for the first time, a provable structure-preserving finite element scheme.
@@ 1.1 Model equations
The model describes the mass transport of the fluid or gas components by convection and diffusion (for the partial mass densities $\rho_{1},\ldots,\rho_{N}$) as well as the momentum balance (for the barycentric velocity $\bm{u}$ and the pressure $p$):
@@
The partial diffusion flux is given as the linear combination of the gradients of the chemical potentials $\mu_{i}$,
@@
where $f$ is the internal energy density and $V_{i}>0$ are the specific volumes of the molecules. A typical example for ideal gas mixtures is given by $f(\vec{\rho})=\sum_{i=1}^{N}\rho_{i}\log(\rho_{i}/\rho)$, used for the numerical tests. Note that our numerical analysis holds for general energy densities that are positively homogeneous of degree one. The incompressibility constraint is imposed by
@@ 1.2 Main results
The discrete energy inequality yields some a priori bounds detailed in Theorem 3.2. However, the existence of a discrete solution to scheme (10)–(13) is still an open problem. The existence of a global weak solution to the continuous problem in a bounded domain with Neumann boundary conditions was shown in Dru21 . The pressure turns out to be a measure with singular values at the minimal and maximal values of the total mass density. Thus, it is not clear to what extend the analysis of Dru21 can be applied to the finite element case. Another approach is the local-in-time existence result of BoDr21 , which yields an integrable pressure function. The idea of the proof is a reformulation of the problem in terms of so-called relative chemical potentials and the construction of a fixed point mapping. As the arguments are quite involved, it is not known whether the arguments hold in the finite-dimensional case.
@@ 1.3 State of the art
The existence of global weak solutions to the incompressible Navier–Stokes–Maxwell–Stefan equations (i.e. $V_{i}=V$ for $i=1,\ldots,N$) was shown in BoPr17 ; ChJu15 ; DoDo21 ; MaTe15 . Existence results for the quasi-incompressible model (i.e. for general $V_{i}>0$) were presented in BoDr21 ; Dru21 . Analytical results for quasi-incompressible fluids, related but different to (7), can be found in, e.g., FLM16 ; GTP15 . In GTP15 , a mass conservative and energy dissipative discontinuous Galerkin finite element discretization for quasi-incompressible Navier–Stokes–Cahn–Hilliard systems was presented. We have only found the work AFMV25 regarding numerical approximations of coupled Maxwell–Stefan fluid models, where a finite element method for coupling the steady-state Maxwell–Stefan equations to compressible Stokes flow was proposed. Thus, structure-preserving numerical methods for quasi-incompressible Navier–Stokes–Maxwell–Stefan systems are missing in the literature.
@@ 2 Various formulations
[…]
@@
    Domain: $\Omega={\mathbb{T}}^{d}$ is the $d$-dimensional torus, $d\geq 1$.
@@
    Mobility matrix: $\mathbb{M}(\vec{\rho})=(M_{ij}(\vec{\rho}))_{i,j=1}^{N}\in(0,\infty)^{N\times N}$ is symmetric positive semi-definite.
@@
Relation (13) follows from the computation in (9), proving the first assertion. If $V_{i}=V$, it follows from $\sum_{i=1}^{N}M_{ij}(\vec{\rho}))=0$ that $\sum_{i=1}^{N}\bm{J}_{i}=0$ and hence ${\operatorname{div}}\bm{u}=0$. Since $\rho=V^{-1}$, we obtain equations (8).
@@
We deduce from (12) and the explicit expression of $f$ that $\rho_{i}/\rho=\exp(\mu_{i}-V_{i}p)$. Then, multiplying this expression by $V_{i}$, summing over $i=1,\ldots,N$, and using the constraint (7), we obtain $01/\rho=\sum_{i=1}^{N}V_{i}\rho_{i}/\rho=\sum_{i=1}^{N}V_{i}\exp(\mu_{i}-V_{i}p)>$ 0. This implies that $\rho>0$ and consequently, $\rho_{i}=\rho\exp(\mu_{i}-V_{i}p)>0$.
@@
00 0 $=\sum_{i=1}^{N}\big{(}\langle\rho_{i}\nabla\mu_{i},\bm{v}\rangle+\langle\mu_{i},{\operatorname{div}}(\rho_{i}\bm{v})\rangle\big{)}$
@@ 3.1 Numerical scheme
We partition the time interval $[0,T]$ uniformly into $n$ subintervals $[t_{k-1},t_{k}]$, where $t_{k}=k\tau$, $k=1,\ldots,n$, and $\tau=T/n$ is the time step size. We introduce
@@
with values in $\phantom{xx}\mbox{with values in }\mathcal{U},$
$\Pi_{\tau}^{1}(\mathcal{U})$ $=\mbox{space of piecewise linear functions on }\{t_{0},\ldots,t_{n}\}\mbox{ with values in }\mathcal{U}.$
@@
$\mathcal{V}_{h}$ =$=\big{\{}v\in H^{1}(\Omega)\cap C^{0}(\overline{\Omega}):v|_{K}\in P_{1}(K)\mbox{ for all }K\in\mathcal{T}_{h}\big{\}},$
$\mathcal{X}_{h}$ =$=\big{\{}v\in H^{1}(\Omega)\cap C^{0}(\overline{\Omega}):v|_{K}\in P_{2}(K)\mbox{ for all }K\in\mathcal{T}_{h}\big{\}},$
@@
Let the initial data $({\vec{\rho}\mkern 2.0mu\vphantom{\rho}}_{h}^{0},\bm{u}_{h}^{0})\in\mathcal{V}_{h,+}^{N}\times\mathcal{X}_{h}^{d}$ be given. The numerical scheme reads as follows: Find $(\vec{\rho}\mkern 2.0mu\vphantom{\rho}_{h}^{k},\bm{u}_{h}^{k})\in\Pi_{\tau}^{1}(\mathcal{V}_{h,+}^{N}\times\mathcal{X}_{h}^{d})$ and $(\vec{\mu}_{h}^{k},p_{h}^{k})\in\Pi_{\tau}^{0}(\mathcal{V}_{h}^{N}\times\mathcal{Q}_{h})$ for $k=0,\ldots,n-1$ such that for $i=1,\ldots,N$,
@@
for all $(\psi_{i},\xi_{i},\bm{v},q)\in\mathcal{V}_{h}^{N}\times\mathcal{V}_{h}^{N}\times\mathcal{X}_{h}^{d}\times\mathcal{Q}_{h}$. Here, $\rho_{h}^{*}$ and $\bm{u}_{h}^{*}$ are functions of $(\rho_{h}^{k},\rho_{h}^{k+1})$ and $(\bm{u}_{h}^{k},\bm{u}_{h}^{k+1})$, respectively. Examples are the explicit scheme $(\rho_{h}^{*},\bm{u}_{h}^{*})=(\rho_{h}^{k},\bm{u}_{h}^{k})$, the implicit scheme $(\rho_{h}^{*},\bm{u}_{h}^{*})=(\rho_{h}^{k+1},\bm{u}_{h}^{k+1})$, or the mid-point rule. In the numerical experiments, we have chosen the explicit approximation. The discrete pressure terms in (22) encode the incompressibility condition (7). In fact, if (7) holds, these terms cancel out.
We mentioned already in the introduction that we assume the existence of a discrete solution to (20)–(23), since a proof is far from being trivial. In particular, we assume that $\rho_{h,i}^{k}$ is nonnegative for $i=1,\ldots,N$ and $k=1,\ldots,n$. This is consistent with the continuous model; see Lemma 2.
@@ 3.2 Properties of the numerical scheme
The following numerical properties hold.
@@
[…]
@@
Set $Z_{h}^{k}:=\sum_{i=1}^{N}V_{i}\rho_{h,i}^{k}-1$. Assumption (A4) implies that $Z_{h}^{0}=0$. By partial mass conservation, $\langle V_{i}\rho_{h,i}^{k},\mathrm{1}\rangle=\langle V_{i}\rho_{h,i}^{0},\mathrm{1}\rangle$ for all $i=1,\ldots,N$ and $k=1,\ldots,n$. Summation over $i=1,\ldots,N$ shows that
@@
Equation (24) writes as
$\langle Z_{h}^{k+1},q\rangle-\langle Z_{h}^{k},q\rangle=\tau\langle Z_{h}^{k}\bm{u}_{h}^{k+1},\nabla q\rangle,$
which reduces with the test function $q=Z_{h}^{1}$ to
$\|q\|_{L^{2}(\Omega)}^{2}=\langle Z_{h}^{1},q\rangle-\langle Z_{h}^{0},q\rangle=\tau\langle Z_{h}^{0}\bm{u}_{h}^{1},\nabla q\rangle=0,$
proving that $q=Z_{h}^{1}=0$ a.e. in $\Omega$. The first assertion follows after an induction argument.
@@ 4 Numerical experiments
In this section, we present some numerical simulations in two space dimensions, illustrating the convergence of the scheme and the structure-preserving properties. The scheme is implemented in NGSolve Sch14 , using Netgen for the generation of an unstructured mesh of triangles. The resulting nonlinear system is solved by Newton’s method with a tolerance of $10^{-9}$ (measured in the $L^{2}(\Omega)$ norm), and the linear systems are computed via direct LU decomposition. In all tests, we fixed the computational domain $\Omega=(0,1)^{2}$ with periodic boundary conditions. This means that we identify $\Omega$ with the two-dimensional torus.
@@ 4.1 Convergence tests
Since no exact solution is available, we compute the error by using the reference solution at the finest space resolution. Given the mesh sizes $h_{k}=2^{-k-1}$ for $k=0,\ldots,7$ and $h_{\rm ref}=2^{-7}$, we define the error quantities
$\mbox{err}(\vec{\rho})=\max_{k=1,\ldots,n}\sum_{i=1}^{2}\|\rho_{h_{k},i}^{k}-\rho_{{\rm ref},i}^{k}\|_{L^{2}(\Omega)}^{2},\quad\mbox{err}(\vec{\mu})=\tau\sum_{k=1}^{n}\sum_{i=1}^{2}\|\mu_{h_{k},i}^{k}-\mu_{{\rm ref},i}^{k}\|_{L^{2}(\Omega)}^{2},$
$\mbox{err}(\bm{u})=\max_{k=1,\ldots,n}\|\bm{u}_{h_{k}}^{k}-\bm{u}_{\rm ref}^{k}\|_{L^{2}(\Omega)}^{2},\qquad\quad\mbox{err}(p)=\tau\sum_{k=1}^{n}\|p_{h_{k}}^{k}-p_{\rm ref}^{k}\|_{L^{2}(\Omega)}^{2}.$
We have chosen the time step size $\tau=10^{-3}$ such that $n=T/\tau=100$. Table 1 presents the errors and the (squared) experimental order of convergence for $V_{1}=0.3$, $V_{2}=0.7$, while Table 2 shows these values for the choice $V_{1}=V_{2}=0.5$. As expected, we observe second-order convergence for $(\vec{\rho},\vec{\mu},p)$ and second-order convergence for $\bm{u}$ (recall that the velocity is approximated by piecewise quadratic polynomials, for which third-order might be possible). In the second test case, the order of convergence for the pressure is larger than in the first test, which comes from the choice $V_{1}=V_{2}$, i.e., we have solved the incompressible Navier–Stokes equations. In both tests, the partial and total masses and the pointwise quasi-incompressibility constraint are conserved up to errors of order $10^{-14}$.
Table 1: $L^{2}(\Omega)$ errors and squared experimental order of convergence (eoc) for $V_{1}=0.3$, $V_{2}=0.7$.
@@
1 $5.31\cdot 10^{-2}$ $3.04\cdot 10^{-3}$ $1.42\cdot 10^{-1}$ $5.67\cdot 10^{-3}$
2 $7.90\cdot 10^{-2}$-0.57 $4.74\cdot 10^{-3}$-0.63 $6.02\cdot 10^{-3}$-4.56 $8.81\cdot 10^{-3}$-0.63
3 $6.83\cdot 10^{-3}$-3.53 $6.15\cdot 10^{-4}$-2.95 $1.48\cdot 10^{-3}$-2.02 $9.94\cdot 10^{-4}$-3.15
4 $3.75\cdot 10^{-4}$-4.18 $4.47\cdot 10^{-5}$-3.78 $1.96\cdot 10^{-4}$-2.90 $6.60\cdot 10^{-5}$-3.91
5 $2.36\cdot 10^{-5}$-4.00 $2.85\cdot 10^{-6}$-3.97 $1.40\cdot 10^{-5}$-3.82 $4.18\cdot 10^{-6}$-3.98
6 $1.42\cdot 10^{-6}$-4.04 $1.69\cdot 10^{-7}$-4.07 $4.66\cdot 10^{-7}$-4.91 $2.49\cdot 10^{-7}$-4.07
7 $9.24\cdot 10^{-8}$-3.94 $1.12\cdot 10^{-8}$-3.91 $1.57\cdot 10^{-8}$-4.89 $1.64\cdot 10^{-8}$-3.92
Table 2: $L^{2}(\Omega)$ errors and squared experimental order of convergence (eoc) for $V_{1}=V_{2}=0.5$.
@@
1 $8.96\cdot 10^{-2}$ $8.14\cdot 10^{-3}$ $1.37\cdot 10^{-1}$ $1.46\cdot 10^{-2}$
2 $1.33\cdot 10^{-1}$-0.57 $6.81\cdot 10^{-3}$-0.26 $4.84\cdot 10^{-4}$-8.15 $1.10\cdot 10^{-2}$-0.42
3 $1.15\cdot 10^{-2}$-3.53 $2.16\cdot 10^{-4}$-4.97 $3.55\cdot 10^{-5}$-3.77 $4.41\cdot 10^{-5}$-7.96
4 $6.36\cdot 10^{-4}$-4.18 $1.71\cdot 10^{-5}$-3.66 $6.26\cdot 10^{-6}$-2.50 $4.78\cdot 10^{-7}$-6.53
5 $3.87\cdot 10^{-5}$-4.04 $1.10\cdot 10^{-6}$-3.96 $4.30\cdot 10^{-7}$-3.86 $1.41\cdot 10^{-8}$-5.09
6 $2.33\cdot 10^{-6}$-4.05 $6.74\cdot 10^{-8}$-4.02 $1.43\cdot 10^{-8}$-4.90 $7.33\cdot 10^{-10}$-4.23
7 $1.56\cdot 10^{-7}$-3.90 $4.51\cdot 10^{-9}$-3.90 $2.89\cdot 10^{-10}$-5.63 $4.80\cdot 10^{-11}$-3.93

### v2

@@ 1 Introduction
Many applications for multicomponent mixtures, like ion transport in biological channels, the dynamics of electrolytes in lithium-ion batteries, or helium–oxygen–carbon dioxide mixtures in the lung, can be described by Maxwell–Stefan equations. In the Fick–Onsager form, they consist of nonlinear parabolic cross-diffusion equations whose diffusion matrix (called mobility matrix) is positive semidefinite only BoDr23 . In the hydrostatic and isothermal case, the variables are the partial mass densities that satisfy a volume-filling constraint implying a constant total mass density. In the general case, the equations are coupled to the compressible or incompressible Navier–Stokes equations. In this paper, we consider the quasi-incompressible model of Dru21 and suggest, for the first time, a provable structure-preserving finite element scheme.
@@ 1.1 Model equations
The model describes the mass transport of the liquid or gas components by convection and diffusion (for the partial mass densities $\rho_{1},\ldots,\rho_{N}$) as well as the momentum balance (for the barycentric velocity $\bm{u}$ and the pressure $p$):
@@
The partial diffusion fluxes are given as linear combinations of the gradients of the chemical potentials $\mu_{i}$,
@@
where $f$ is the internal energy density and $V_{i}>0$ are the constant specific volumes of the molecules. A typical example for ideal gas mixtures is given by $f(\vec{\rho})=\sum_{i=1}^{N}\rho_{i}\log(\rho_{i}/\rho)$, used for the numerical tests. Note that our numerical analysis holds for general energy densities that are positively homogeneous of degree one. The incompressibility constraint is imposed by
@@
An analogous model can be derived when the quasi-incompressibility constraint (7) is neglected. In this case, no equation for the divergence (13) and no contribution of the pressure $p$ in the chemical potentials appear, cf. Dru21b ; Eike25 . Therefore, the system needs to be considered with an equation of state for the pressure.
@@ 1.2 Main results
The discrete energy inequality yields some a priori bounds detailed in Theorem 3.2. However, the existence of a discrete solution to scheme (10)–(13) is still an open problem. The existence of a global weak solution to the continuous problem in a bounded domain with Neumann boundary conditions was shown in Dru21 . The pressure turns out to be a measure with singular values at the minimal and maximal values of the total mass density. Thus, it is not clear to what extent the analysis of Dru21 can be applied to the finite element case. Another approach is the local-in-time existence result of BoDr21 , which yields an integrable pressure function. The idea of the proof is a reformulation of the problem in terms of so-called relative chemical potentials and the construction of a fixed point mapping. As the arguments are quite involved, it is not known whether the arguments hold in the finite-dimensional case.
@@ 1.3 State of the art
The existence of global weak solutions to the incompressible Navier–Stokes–Maxwell–Stefan equations (i.e. $V_{i}=V$ for $i=1,\ldots,N$) was shown in BoPr17 ; ChJu15 ; DoDo21 ; MaTe15 . Existence results for the quasi-incompressible model (i.e. for general $V_{i}>0$) were presented in BoDr21 ; Dru21 . Analytical results for quasi-incompressible fluids, related but different to (7), can be found in, e.g., FLM16 ; GTP15 . In GTP15 , a mass conservative and energy dissipative discontinuous Galerkin finite element discretization for quasi-incompressible Navier–Stokes–Cahn–Hilliard systems was presented. In AFMV25 ; Van-Brunt22 numerical approximations of coupled Maxwell-–Stefan fluid models using finite elements methods are considered, where steady-state Maxwell-–Stefan equations are coupled with a compressible Stokes flow. Similar steady-state models and discretizations in the quasi-incompressible case are studied in Baier25 , using an explicit formulation of the forces instead of the fluxes. To include additional effects, like in VBru23 , the explicit force formulation might be more appropriate. However, structure-preserving numerical methods for non-stationary quasi-incompressible Navier–Stokes–Maxwell–Stefan systems are missing in the literature.
@@ 2 Various formulations
We introduce the following notation. Vectors in ${\mathbb{R}}^{N}$ are written as $\vec{v}$, while vectors in ${\mathbb{R}}^{d}$ are denoted by $\bm{v}$. Matrices are formulated as $\mathbb{M}$. The $L^{2}(\Omega)$ inner product is denoted by $\langle\cdot,\cdot\rangle$, and we set $\mathrm{1}=(1,\ldots,1)^{T}\in{\mathbb{R}}^{N}$. We impose the following assumptions:
@@
    Domain: let $\Omega=(0,1)^{d}$ be topologically equivalent to the $d$-dimensional torus, $d\geq 1$. This means we enforce periodic boundary conditions.
@@
    Mobility matrix: $\mathbb{M}(\vec{\rho})=(M_{ij}(\vec{\rho}))_{i,j=1}^{N}\in(0,\infty)^{N\times N}$ is symmetric positive semi-definite with $\mathrm{1}\in\mathrm{ker}(\mathbb{M})$.
@@
Relation (13) follows from the computation in (9), proving the first assertion. If $V_{i}=V$, it follows from $\sum_{i=1}^{N}M_{ij}(\vec{\rho}))=0$, i.e. (A1), that $\sum_{i=1}^{N}\bm{J}_{i}=0$ and hence ${\operatorname{div}}\bm{u}=0$. Since $\rho=V^{-1}$, we obtain equations (8).
@@
We deduce from (12) and the explicit expression of $f$ that $\rho_{i}/\rho=\exp(\mu_{i}-V_{i}p)$. Then, multiplying this expression by $V_{i}$, summing over $i=1,\ldots,N$, and using the constraint (7), we obtain $1/\rho=\sum_{i=1}^{N}V_{i}\rho_{i}/\rho=\sum_{i=1}^{N}V_{i}\exp(\mu_{i}-V_{i}p)>0$. This implies that $\rho>0$ and consequently, $\rho_{i}=\rho\exp(\mu_{i}-V_{i}p)>0$.
@@
To motivate our structure-preserving numerical method, we first derive a suitable weak formulation of (1)–(4), (6)–(7), such that the relevant structures as conservation of partial masses and the energy dissipation can be achieved by using suitable test functions. For instance, to derive the contribution of the kinetic energy from the standard conservative variables, we need to test the momentum equation with $\bm{u}$ and the total mass equation with $\frac{1}{2}|\bm{u}|^{2}$. The approximate velocity is typically contained in an approximation space, while $\frac{1}{2}|\bm{u}|^{2}$ is usually not in the approximation space for the densities.
@@
$0$ $=\sum_{i=1}^{N}\big(\langle\rho_{i}\nabla\mu_{i},\bm{v}\rangle+\langle\mu_{i},{\operatorname{div}}(\rho_{i}\bm{v})\rangle\big)$
@@ 3.1 Numerical scheme
We partition the time interval $[0,T]$ uniformly into $n$ subintervals $t_{k-1},t_{k}]$, where $t_{k}=k\tau$, $k=1,\ldots,n$, and $\tau=T/n$ is the time step size. We introduce
$\Pi_{\tau}^{0}(\mathcal{U})$ $=\mbox{space of discontinuous piecewise constant functions on }\{t_{0},\ldots,t_{n}\}$
$\phantom{xx}\mbox{with values in }\mathcal{U},$
@@
$\mathcal{V}_{h}$ $=\big\{v\in H^{1}(\Omega)\cap C^{0}(\overline{\Omega}):v|_{K}\in P_{1}(K)\mbox{ for all }K\in\mathcal{T}_{h}\big\},$
$\mathcal{X}_{h}$ $=\big\{v\in H^{1}(\Omega)\cap C^{0}(\overline{\Omega}):v|_{K}\in P_{2}(K)\mbox{ for all }K\in\mathcal{T}_{h}\big\},$
@@
Let the initial data $({\vec{\rho}\mkern 2.0mu\vphantom{\rho}}_{h}^{0},\bm{u}_{h}^{0})\in\mathcal{V}_{h}^{N}\times\mathcal{X}_{h}^{d}$ be given. The numerical scheme reads as follows: Find $(\vec{\rho}\mkern 2.0mu\vphantom{\rho}_{h}^{k},\bm{u}_{h}^{k})\in\Pi_{\tau}^{1}(\mathcal{V}_{h}^{N}\times\mathcal{X}_{h}^{d})$ and $(\vec{\mu}_{h}^{k},p_{h}^{k})\in\Pi_{\tau}^{0}(\mathcal{V}_{h}^{N}\times\mathcal{Q}_{h})$ for $k=0,\ldots,n-1$ such that for $i=1,\ldots,N$,
@@
for all $(\psi_{i},\xi_{i},\bm{v},q)\in\mathcal{V}_{h}^{N}\times\mathcal{V}_{h}^{N}\times\mathcal{X}_{h}^{d}\times\mathcal{Q}_{h}$. Here, $\rho_{h}^{*}$ and $\bm{u}_{h}^{*}$ are functions of $(\rho_{h}^{k},\rho_{h}^{k+1})$ and $(\bm{u}_{h}^{k},\bm{u}_{h}^{k+1})$, respectively. Examples are the explicit scheme $(\rho_{h}^{*},\bm{u}_{h}^{*})=(\rho_{h}^{k},\bm{u}_{h}^{k})$, the implicit scheme $(\rho_{h}^{*},\bm{u}_{h}^{*})=(\rho_{h}^{k+1},\bm{u}_{h}^{k+1})$, or the mid-point rule. In the numerical experiments, we have chosen the explicit approximation. The discrete pressure terms in (22) encode the incompressibility condition (7). In fact, if (7) holds, these terms cancel out. This means that if the discrete solution satisfies the constraint (7), the discrete pressure does not appear in (22) directly, but only implicitly via the chemical potentials.
We mentioned already in the introduction that we assume the existence of a discrete solution to (20)–(23), since a proof is far from being trivial. In particular, we expect that $\rho_{h,i}^{k}$ is nonnegative for $i=1,\ldots,N$ and $k=1,\ldots,n$. This is consistent with the continuous model; see Lemma 2. Hence, we expect that the positivity of the discrete partial densities is a consequence of the singular behavior of the free energy density or its derivatives.
@@ 3.2 Properties of the numerical scheme
The following numerical properties hold, given that the discrete solutions are positive.
@@
$\frac{1}{\tau}\langle E^{k+1}_{h}-E^{k}_{h},1\rangle$ $=-\mathcal{D}_{\rm num}-\big\langle\mathbb{S}(\vec{\rho}\mkern 2.0mu\vphantom{\rho}_{h}^{*},\nabla\bm{u}_{h}^{k+1}),\nabla\bm{u}_{h}^{k+1}\big\rangle$ (29)
@@
Set $Z_{h}^{k}:=\sum_{i=1}^{N}V_{i}\rho_{h,i}^{k}-1$. Assumption (A4) implies that $Z_{h}^{0}=0$. We use an induction argument to show that $Z_{h}^{k}=0$ for all $k$. We wish to take $Z_{h}^{k+1}$ as a test function in (24). To this end, we first show that $Z_{h}^{k+1}\in\mathcal{Q}_{h}$. By partial mass conservation, $\langle V_{i}\rho_{h,i}^{k},\mathrm{1}\rangle=\langle V_{i}\rho_{h,i}^{0},\mathrm{1}\rangle$ for all $i=1,\ldots,N$ and $k=1,\ldots,n$. Summation over $i=1,\ldots,N$ shows that
@@
Equation (24) with the test function $q=Z_{h}^{k+1}\in\mathcal{Q}_{h}$ reads as
$\|Z_{h}^{k+1}\|_{L^{2}(\Omega)}^{2}+\|Z_{h}^{k+1}-Z_{h}^{k}\|_{L^{2}(\Omega)}^{2}=\|Z_{h}^{k}\|_{L^{2}(\Omega)}^{2}+\tau\langle Z_{h}^{k}\bm{u}_{h}^{k+1},\nabla Z_{h}^{k+1}\rangle=0.$
We conclude from the above equation that if $Z_{h}^{k}=0$ everywhere, then also $Z_{h}^{k+1}=0$ everywhere. This shows the claim.
@@ 4 Numerical experiments
In this section, we present some numerical simulations in two space dimensions, illustrating the convergence of the scheme and the structure-preserving properties. The scheme is implemented in NGSolve Sch14 , using Netgen for the generation of an unstructured mesh of triangles. The resulting nonlinear system is solved by Newton’s method with a tolerance of $10^{-9}$ (measured in the $L^{2}(\Omega)$ norm of the Newton updates), and the linear systems are computed via direct LU decomposition. In all tests, we fixed the computational domain $\Omega=(0,1)^{2}$ with periodic boundary conditions. This means that we identify $\Omega$ with the two-dimensional torus.
@@ 4.1 Convergence tests
Since no exact solution is available, we compute the error by using the reference solution at the finest space resolution. Given the mesh sizes $h_{k}=2^{-k-1}$ for $k=1,\ldots,8$ and $h_{\rm ref}=2^{-8}$, we define the error quantities
$\mbox{err}(\vec{\rho})=\max_{\ell=1,\ldots,n}\sqrt{\sum_{i=1}^{2}\|\rho_{h_{k},i}^{\ell}-\rho_{{\rm ref},i}^{\ell}\|_{L^{2}(\Omega)}^{2}},\quad\mbox{err}(\vec{\mu})=\sqrt{\tau\sum_{\ell=1}^{n}\sum_{i=1}^{2}\|\mu_{h_{k},i}^{\ell}-\mu_{{\rm ref},i}^{\ell}\|_{L^{2}(\Omega)}^{2}},$
$\mbox{err}(\bm{u})=\max_{\ell=1,\ldots,n}\sqrt{\|\bm{u}_{h_{k}}^{\ell}-\bm{u}_{\rm ref}^{\ell}\|_{L^{2}(\Omega)}^{2}},\qquad\quad\mbox{err}(p)=\sqrt{\tau\sum_{\ell=1}^{n}\|p_{h_{k}}^{\ell}-p_{\rm ref}^{\ell}\|_{L^{2}(\Omega)}^{2}}.$
We have chosen the time step size $\tau=10^{-3}$ such that $n=T/\tau=100$. Table 1 presents the errors and the experimental order of convergence for $V_{1}=0.3$, $V_{2}=0.7$, while Table 2 shows these values for the choice $V_{1}=V_{2}=0.5$. As expected, we observe second-order convergence for $(\vec{\rho},\vec{\mu},p)$ and second-order convergence for $\bm{u}$ (recall that the velocity is approximated by piecewise quadratic polynomials, for which third-order might be possible). In the second test case, the order of convergence for the velocity is larger than in the first test, which comes from the choice $V_{1}=V_{2}$, i.e., we have solved the incompressible Navier–Stokes equations. In fact, for the incompressible Navier–Stokes system, one can show optimal third-order convergence, while in the quasi-incompressible case, it is unclear if that order can be proven. In both tests, the partial and total masses and the point wise quasi-incompressibility constraint are conserved up to errors of order $10^{-14}$. The time step is chosen sufficiently small for two reasons. The first is to ensure that the temporal error does not dominate the spatial error. The second is to guarantee convergence of Newton iterations for all meshes. The latter is related to the existence of (unique) discrete solutions.
Table 1: $L^{2}(\Omega)$ errors and experimental order of convergence for $V_{1}=0.3$, $V_{2}=0.7$.
@@
1 $2.30\cdot 10^{-1}$ $5.52\cdot 10^{-2}$ $3.77\cdot 10^{-1}$ $7.53\cdot 10^{-2}$
2 $2.81\cdot 10^{-1}$ -0.29 $6.89\cdot 10^{-2}$ -0.32 $7.76\cdot 10^{-2}$ -2.28 $9.39\cdot 10^{-2}$ -0.32
3 $8.27\cdot 10^{-2}$ -1.77 $2.48\cdot 10^{-2}$ -1.48 $3.85\cdot 10^{-2}$ -1.01 $3.15\cdot 10^{-2}$ -1.58
4 $1.94\cdot 10^{-2}$ -2.09 $6.68\cdot 10^{-3}$ -1.89 $1.40\cdot 10^{-2}$ -1.45 $8.12\cdot 10^{-3}$ -1.96
5 $4.86\cdot 10^{-3}$ -2.00 $1.69\cdot 10^{-3}$ -1.99 $3.74\cdot 10^{-3}$ -1.91 $2.04\cdot 10^{-3}$ -1.99
6 $1.19\cdot 10^{-3}$ -2.02 $4.11\cdot 10^{-4}$ -2.04 $6.82\cdot 10^{-4}$ -2.46 $4.99\cdot 10^{-4}$ -2.04
7 $3.04\cdot 10^{-4}$ -1.97 $1.06\cdot 10^{-4}$ -1.96 $1.25\cdot 10^{-4}$ -2.45 $1.28\cdot 10^{-4}$ -1.96
Table 2: $L^{2}(\Omega)$ errors and experimental order of convergence for $V_{1}=V_{2}=0.5$.
@@
1 $2.99\cdot 10^{-1}$ $9.02\cdot 10^{-2}$ $3.70\cdot 10^{-1}$ $1.21\cdot 10^{-1}$
2 $3.64\cdot 10^{-1}$ -0.29 $8.26\cdot 10^{-2}$ -0.13 $2.20\cdot 10^{-2}$ -4.08 $1.05\cdot 10^{-1}$ -0.21
3 $1.07\cdot 10^{-1}$ -1.77 $1.47\cdot 10^{-2}$ -2.49 $5.96\cdot 10^{-3}$ -1.89 $6.64\cdot 10^{-3}$ -3.98
4 $2.52\cdot 10^{-2}$ -2.09 $4.14\cdot 10^{-3}$ -1.83 $2.50\cdot 10^{-3}$ -1.25 $6.91\cdot 10^{-4}$ -3.27
5 $6.22\cdot 10^{-3}$ -2.02 $1.05\cdot 10^{-3}$ -1.98 $6.56\cdot 10^{-4}$ -1.93 $1.19\cdot 10^{-4}$ -2.55
6 $1.53\cdot 10^{-3}$ -2.02 $2.60\cdot 10^{-4}$ -2.01 $1.20\cdot 10^{-4}$ -2.45 $2.71\cdot 10^{-5}$ -2.12
7 $3.95\cdot 10^{-4}$ -1.95 $6.72\cdot 10^{-5}$ -1.95 $1.70\cdot 10^{-5}$ -2.82 $6.93\cdot 10^{-6}$ -1.97